OpenStudy (mpj4):

can x/(x^2 + 1) be simplified to 1/x + x?

OpenStudy (mpj4):

\[\frac{x}{x^2+1} = \frac{1}{x} + x\]

rvc (rvc):

nope

OpenStudy (mpj4):

ah, no wonder I got my answer wrong. Thanks